OpenStudy (anonymous):

The Good Seed Flower Shop sells its own brand of fertilizer. It mixes one brand of fertilizer that contains 7% nitrogen with another than contains 8% nitrogen. The resulting mixture contains 7.7% nitrogen. How many pounds of the 8% fertilizer must be used to make 200 pounds of the 7.7% mixture?

OpenStudy (nurali):

.08X+.07(200-X)=.077*200 .08X+14-.07X=15.4 .01X=15.4-14 .01X=1.4 X=1.4/.01 X=140 POUNDS OF 8% FERTILIZER IS NEEDED. 200-140=60 POUNDS OF 7% IS NEEDED. PROOF: .08*140+.07*60=.077*200 11.2+4.2=15.4 15.4=15.4